Find an equation of the tangent line to the curve x⁴ + 2y⁴ = 33 at the point (1,2).

1
Expert's answer
2021-12-06T16:43:41-0500
"x^4+2y^4=33"

Differentiate both sides with respect to "x"


"4x^3+8y^3y'=0"

"y'=-\dfrac{x^3}{2y^3}"

Point (1,2)

"slope=m=-\dfrac{1^3}{2(2^3)}=-\dfrac{1}{16}"

The equation of the tangent line in point-slope form


"y-2=-\dfrac{1}{16}(x-1)"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{1}{16}x+\dfrac{33}{16}"